# Implementing Grassroots Logic Programs with Multiagent Transition Systems and AI

### v1

@@ D.1 Protocols and the Grassroots Property
A local-states function $S:2^{\Pi}\to 2^{\mathcal{S}}$ maps every finite set of agents $P\subset\Pi$ to a set of local states $S(P)\subseteq\mathcal{S}$ that includes $s0$ and satisfies $P\subset P^{\prime}\implies S(P)\subseteq S(P^{\prime})$.
@@
Let $\emptyset\subset P\subset P^{\prime}\subset\Pi$. If $c^{\prime}$ is a configuration over $P^{\prime}$, then $c^{\prime}/P$, the projection of $c^{\prime}$ onto $P$, is the configuration $c$ over $P$ defined by $c_{p}:=c^{\prime}_{p}$ for every $p\in P$.
@@
    oblivious if for every $\emptyset\subset P\subset P^{\prime}\subset\Pi$, $T(P){\uparrow}P^{\prime}\subseteq T(P^{\prime})$
@@
    interactive if for every $\emptyset\subset P\subset P^{\prime}\subset\Pi$ and every configuration $c\in C(P^{\prime})$ such that $c/P\in C(P)$, there is a computation $c\xrightarrow{*}c^{\prime}$ of $\mathcal{F}(P^{\prime})$ for which $c^{\prime}/P\notin C(P)$

### v2

@@ D.1 Protocols and the Grassroots Property
A local-states function $S:2^{\Pi}\to 2^{\mathcal{S}}$ maps every finite set of agents $P\subset\Pi$ to a set of local states $S(P)\subset\mathcal{S}$ that includes $s0$ and satisfies $P\subset P^{\prime}\implies S(P)\subset S(P^{\prime})$.
@@
Let $\emptyset\subset P\subset P^{\prime}\subset\Pi$. If $c^{\prime}$ is a configuration over $P^{\prime}$, then $c^{\prime}/P$, the projection of $c^{\prime}$ over $P$, is the configuration $c$ over $P$ defined by $c_{p}:=c^{\prime}_{p}$ for every $p\in P$.
@@
    oblivious if for every $\emptyset\subset P\subset P^{\prime}\subseteq\Pi$, $T(P){\uparrow}P^{\prime}\subseteq T(P^{\prime})$
@@
    interactive if for every $\emptyset\subset P\subset P^{\prime}\subseteq\Pi$ and every configuration $c\in C(P^{\prime})$ such that $c/P\in C(P)$, there is a computation $c\xrightarrow{*}c^{\prime}$ of $\mathcal{F}(P^{\prime})$ for which $c^{\prime}/P\notin C(P)$